An inexpensive and accurate method of measuring the quantity of electricity passing through a circuit is to pass it through a solution of a metal ion and weigh the metal deposited. A silver electrode immersed in an Ag+ solution weighs 1.7854 g before the current passes and weighs 1.9283 g after the current has passed.

How many coulombs have passed?
Answer #1

After the current passed the solution weighs 1.9283 and before it weighs 1.7854..the amount of silver deposited on electrode is 1.9283-1.7854= 0.1429 gms

the molar weight of silver is 108gms

the no.of moles of silevr deposited will be 0.1429/108= 0.0013 moles

if we pass 1faraday of current 1 mole will be deposited...but for 0.0013 moles 0.0013 Faraday of current is required...

1 Faraday is equal to96485 coloumbs

so, 0.0013*96485= 125.4305 coloumbs of current should be passed